# Mean-field underdamped Langevin dynamics and its space-time discretization

### 01 Jan 2024 v2

@@ 4.1 Training mean-field neural networks
Consider a two-layer neural network in the mean-field regime (with infinite neurons), which can be parametrized as $h(\mu;a)\triangleq\mathbb{E}_{\mu}[h(x;a)],$ where $h(x;a)$ represents a single neuron with trainable parameter $x$ and input $a$ (e.g. $h(x;a)=\sigma(x^{\mathsf{T}}a)$ for activation function $\sigma$); $\mu$ is the probability distribution of the parameter $x$. Given dataset $(a_{i},b_{i})_{i=1}^{n}$, loss function $l$ and regularization function $r$, we can choose $F(\mu)$ in (1) to be
@@
Since $l$ is convex and $h(\mu;a)$ is a linear functional of $\mu$, which is also convex, their composition remains convex if $h$ satisfies quadratic growth (Proposition 3.2 in Chen et al. (2022)). The learning of two-layer neural networks is thus transformed into a convex optimization problem in the space of measures, for which the convexity of loss function can be utilized to show the global convergence of the gradient-based optimization methods (Nitanda and Suzuki, 2017; Mei et al., 2018; Chizat and Bach, 2018; Rotskoff and Vanden-Eijnden, 2018b; Sirignano and Spiliopoulos, 2020; Suzuki et al., 2023). The entropy regularization term in (1) comes from the noisy gradient descent of $F(\mu)$. While the minimizers of $F(\mu)+\lambda\text{Ent}(\mu)$ and $F(\mu)$ are different, the global convergence analysis of $F(\mu)+\lambda\text{Ent}(\mu)$ requires less stringent assumptions (Rotskoff and Vanden-Eijnden, 2018a; Mei et al., 2019; Hu et al., 2019; Nitanda et al., 2022) and we can control the value of $\lambda$ to make the minimizers of $F(\mu)+\lambda\text{Ent}(\mu)$ close to the minimizers of $F(\mu)$.
@@ 4.2 Density estimation
where $\lambda^{\prime}>0$ and $r$ is the regularization function. Suzuki et al. (2023) show that objective (15) satisfies Assumptions 2.1-2.4 by choosing smooth and light-tailed kernel $k$, such as Gaussian radial basis function (RBF) kernel defined as $k(z,z^{\prime})\triangleq\exp(-\|z-z^{\prime}\|^{2}/2\sigma^{\prime 2})$ for $\sigma^{\prime}>0$. With Gaussian RBF kernel $k$ ($\sigma^{\prime}=\sigma$), we demonstrate that objective (15) also satisfies our Assumptions 2.5-2.7 if we choose $\lambda^{\prime}\lesssim 2\pi$, $r(x)=\|x\|^{2}/2$ and $\mu_{0}=\mathcal{N}(0,I_{2d})$. We defer the verification of our assumptions in the setting above to Appendix B.
@@ 5 Numerical experiments
Figure 2: UNLA with different number of particles
@@ Appendix A Helpful lemmas
$\leq h\|v_{kh}\|+\left\|\int_{kh}^{t}\int_{0}^{\tau}\gamma v_{\tau^{\prime}}\mathrm{d}\tau^{\prime}\mathrm{d}\tau\right\|+\left\|\int_{kh}^{t}\int_{kh}^{\tau}{D}_{\mu}F(\mu_{\tau^{\prime}}^{x},x_{\tau^{\prime}})\mathrm{d}\tau^{\prime}\mathrm{d}\tau\right\|+\left\|\int_{kh}^{t}\int_{kh}^{\tau}\sqrt{2\gamma}\mathrm{d}\mathrm{B}_{\tau^{\prime}}\mathrm{d}\tau\right\|$
@@
$\quad+\left\|\int_{kh}^{t}\int_{kh}^{\tau}\sqrt{2\gamma}\mathrm{d}\mathrm{B}_{\tau^{\prime}}\mathrm{d}\tau\right\|$
@@
$\|v_{t}-v_{kh}\|$ $=\left\|\int_{kh}^{t}\gamma v_{\tau}\mathrm{d}\tau\right\|+\left\|\int_{kh}^{t}{D}_{\mu}F(\mu_{\tau}^{x},x_{\tau})\mathrm{d}\tau\right\|+\left\|\int_{kh}^{t}\sqrt{2\gamma}\mathrm{d}\mathrm{B}_{t}\right\|$
@@
$|\text{Ent}(\mu_{0})|=\left|\int\mu_{0}\log\mu_{0}\right|=\frac{d}{2}\log(2\pi)+\frac{1}{2}\mathbb{E}_{\mu_{0}}\|\cdot\|^{2}\lesssim d.$
@@
$\mathbb{E}\|(x_{t},v_{t})\|^{2}\leq\frac{2}{\mathscr{C}_{\textsf{LSI}}}\mathcal{F}(\mu_{0})+2\textbf{m}_{2}^{2}\lesssim\frac{\mathscr{L}d}{\mathscr{C}_{\textsf{LSI}}}+d$
@@
$\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\|(x^{i}_{t},v^{i}_{t})\|^{2}=\frac{1}{N}\sum_{i=1}^{N}W_{2}^{2}(\mu^{i}_{t},\delta_{0})$ $\leq\frac{2}{N}\sum_{i=1}^{N}W_{2}^{2}(\mu^{i}_{t},\mu_{*})+2W_{2}^{2}(\mu_{*},\delta_{0})$
@@
The second inequality follows from Talagrand’s inequality which can be implied by Assumption 2.3. The third inequality follows from Lemma 4. The fourth inequality follows from Lemma 3 and the last inequality follows that $\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{F}^{N}(\mu^{N}_{t})<0$ along the NULD (Proof of Theorem 2.2 in Chen et al. (2023)) and $\mathcal{F}(\mu_{*})\geq 0$. By the definition of $\mathcal{F}^{N}(\mu^{N})$, we have $\mathcal{F}^{N}(\mu_{0}^{N})=\int(NF(\mu_{\textbf{x}})+\frac{1}{2}\|\textbf{v}\|^{2})\mu_{0}^{N}(\mathrm{d}\textbf{x}\mathrm{d}\textbf{v})+\text{Ent}(\mu_{0}^{N})$. Similar to the proof of Lemma 7, since $(\textbf{x},\textbf{v})\sim\mathcal{N}^{\otimes N}(0,I_{2d})$, we have $\int\frac{1}{2}\|\textbf{v}\|^{2}\mu_{0}^{N}(\mathrm{d}\textbf{x}\mathrm{d}\textbf{v})\lesssim Nd$ and $|\text{Ent}(\mu_{0}^{N})|\lesssim Nd$. By Assumption 2.7 and Assumption 2.5, we also have $\int NF(\mu_{\textbf{x}})\mu_{0}^{N}(\mathrm{d}\textbf{x}\mathrm{d}\textbf{v})\lesssim N\mathscr{L}d$ and $\textbf{m}_{2}^{2}\lesssim d$. Thus we have
@@
$\mathrm{d}x_{t}$ $=b_{t}^{\textbf{P}}\mathrm{d}t+\sigma\mathrm{d}\mathrm{B}^{\textbf{P}}_{t}\quad\text{under }\textbf{P}_{T},$
$\mathrm{d}x_{t}$ $=b_{t}^{\textbf{Q}}\mathrm{d}t+\sigma\mathrm{d}\mathrm{B}^{\textbf{Q}}_{t}\quad\text{under }\textbf{Q}_{T},$
@@
$\frac{\mathrm{d}\mathbf{P}_{T}}{\mathrm{d}\mathbf{Q}_{T}}=\exp\left(\int_{0}^{T}\langle\sigma_{t}^{{\dagger}}(b_{t}^{\textbf{P}_{T}}-b_{t}^{\textbf{Q}_{T}}),\mathrm{d}\mathrm{B}^{\textbf{Q}_{T}}_{t}\rangle-\frac{1}{2}\int_{0}^{T}\|\sigma_{t}^{{\dagger}}(b_{t}^{\textbf{P}_{T}}-b_{t}^{\textbf{Q}_{T}})\|^{2}\mathrm{d}t\right)$
Besides, $(\tilde{\mathrm{B}}_{t})_{t\in[0,T]}$ defined by $\mathrm{d}\tilde{\mathrm{B}}_{t}\triangleq\mathrm{d}\mathrm{B}_{t}+\sigma_{t}^{{\dagger}}(b_{t}^{Y}-b_{t}^{X})$ is a $\mathbf{P}_{T}$-Brownian motion.
@@ Appendix B Verification of assumptions
B.2 Verification of Assumption 2.6
@@
  • If $|h(x;a)|\leq\sqrt{\mathscr{L}}(1+\|x\|)$, we have $|h(\mu_{0};a)|\leq\sqrt{\mathscr{L}}\int(1+\|x\|)\mu_{0}(dx)\lesssim\sqrt{\mathscr{L}}d^{1/2}$. If $l$ is $\sqrt{\mathscr{L}}$-Lipschitz, we have $|l(h(\mu_{0};a_{i}),b_{i})|\leq\sqrt{\mathscr{L}}|h(\mu_{0};a_{i})-b_{i}|\lesssim\mathscr{L}d^{1/2}+\sqrt{\mathscr{L}}\max_{i=1}^{n}|b_{i}|$. We can normalize the data samples to ensure $\max_{i=1}^{n}|b_{i}|\lesssim d\land\sqrt{\mathscr{L}}$. Thus we have $F(\mu_{0})=O(\mathscr{L}d+d)$.
Density estimation
@@
B.3 Verification of Assumption 2.7

### 17 Jan 2024 v3

@@ 4.1 Training mean-field neural networks
Consider a two-layer neural network in the mean-field regime (with infinite neurons), which can be parametrized as $h(\mu;a)\triangleq\mathbb{E}_{\mu}[h(x;a)],$ where $h(x;a)$ represents a single neuron with trainable parameter $x$ and input $a$ (e.g. $h(x;a)=\sigma(x^{\mathsf{T}}a)$ for activation function $\sigma$); $\mu$ is the probability distribution of the parameter $x$. Given dataset $(a_{i},b_{i})_{i=1}^{n}$, loss function $l$ and regularization function $r$, we can choose $F(\mu)$ in (2.2) to be
@@
Since $l$ is convex and $h(\mu;a)$ is a linear functional of $\mu$, which is also convex, their composition remains convex if $h$ satisfies quadratic growth (Proposition 3.2 in Chen et al. (2022)). The learning of two-layer neural networks is thus transformed into a convex optimization problem in the space of measures, for which the convexity of loss function can be utilized to show the global convergence of the gradient-based optimization methods (Nitanda and Suzuki, 2017; Mei et al., 2018; Chizat and Bach, 2018; Rotskoff and Vanden-Eijnden, 2018b; Sirignano and Spiliopoulos, 2020; Suzuki et al., 2023). The entropy regularization term in (2.2) comes from the noisy gradient descent of $F(\mu)$. While the minimizers of $F(\mu)+\lambda\text{Ent}(\mu)$ and $F(\mu)$ are different, the global convergence analysis of $F(\mu)+\lambda\text{Ent}(\mu)$ requires less stringent assumptions (Rotskoff and Vanden-Eijnden, 2018a; Mei et al., 2019; Hu et al., 2019; Nitanda et al., 2022) and we can control the value of $\lambda$ to make the minimizers of $F(\mu)+\lambda\text{Ent}(\mu)$ close to the minimizers of $F(\mu)$.
@@ 4.2 Density estimation via MMD minimization
where $\lambda^{\prime}>0$ and $r$ is the regularization function. Suzuki et al. (2023) show that objective (15) satisfies Assumptions 2.1, 2.3 and 2.4 by choosing smooth and light-tailed kernel $k$, such as Gaussian radial basis function (RBF) kernel defined as $k(z,z^{\prime})\triangleq\exp(-\|z-z^{\prime}\|^{2}/2\sigma^{\prime 2})$ for $\sigma^{\prime}>0$. We also verify that objective (15) also satisfies our Assumption 2.2 with the same choice of kernel. With Gaussian RBF kernel $k$ ($\sigma^{\prime}=\sigma$), we demonstrate that objective (15) satisfies Assumptions 2.5-2.7 if we specify $\lambda^{\prime}\lesssim 2\pi$, $r(x)=\|x\|^{2}/2$, $\mu_{0}=\mathcal{N}(0,I_{2d})$ and $\mu_{0}^{N}=\mathcal{N}(0,I_{2Nd})$.
4.3 Kernel Stein discrepancy minimization
Kernel Stein discrepancy (KSD) minimization is a method for sampling from a target distribution $\mu_{*}$ if we have the access to the score function $s_{\mu_{*}}(x)=\nabla\log\mu_{*}(x)$ (Chwialkowski et al., 2016; Liu et al., 2016). For example, if $\mu_{*}(x)\propto\exp(-V(x))$, we need the access of $\nabla V(x)$. For a positive definite kernel $k$, the Stein kernel is defined as
$u_{\mu_{*}}(x,x^{\prime})=s_{\mu_{*}}^{\mathsf{T}}(x)k(x,x^{\prime})s_{\mu_{*}}(x^{\prime})+s_{\mu_{*}}^{\mathsf{T}}(x)\nabla_{x^{\prime}}k(x,x^{\prime})+\nabla_{x}^{\mathsf{T}}k(x,x^{\prime})s_{\mu_{*}}(x^{\prime})+\textsf{tr}(\nabla_{x,x^{\prime}}k(x,x^{\prime})).$
The KSD between $\mu$ and $\mu_{*}$ is defined as
$\textsf{KSD}(\mu)=\iint u_{\mu_{*}}(x,x^{\prime})\mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime}).$ (16)
We can choose $F(\mu)$ in (2.2) to be
$F(\mu)=\textsf{KSD}(\mu)+\lambda^{\prime}\mathbb{E}_{x\sim\mu}[r(x)],$ (17)
where $\lambda^{\prime}>0$ and $r$ is the regularization function. Suzuki et al. (2023) show that objective (17) satisfies Assumptions 2.1, 2.3 and 2.4 by choosing light-tailed kernel and assume the score function satisfies
$\max\{\|\nabla\log\mu_{*}(x)\|,\|\nabla^{\otimes 2}\log\mu_{*}(x)\|_{\textsf{op}},\|\nabla^{\otimes 3}\log\mu_{*}(x)\|_{\textsf{op}}\}\leq\mathscr{L}(1+\|x\|).$ (18)
More specifically, if $\mu_{*}\propto\exp(-V)$, the potential function $V$ should satisfies
$\max\{\|\nabla V(x)\|,\|\nabla^{\otimes 2}V(x)\|_{\textsf{op}},\|\nabla^{\otimes 3}V(x)\|_{\textsf{op}}\}\leq\mathscr{L}(1+\|x\|),$
which subsumes many distributions. With the same choice of kernel in Suzuki et al. (2023), we verify that objective (17) also satisfies Assumption 2.2 and satisfies our Assumptions 2.5-2.7 with $\lambda^{\prime}\lesssim(2\pi)^{\frac{d}{d-2}}\exp\left(-\frac{4\mathscr{L}}{d-2}\right)$, $r(x)=\|x\|^{2}/2$, $\mu_{0}=\mathcal{N}(0,I_{d})$ and $\mu_{0}^{N}=\mathcal{N}(0,I_{2Nd})$.
@@ 5 Numerical experiments
Figure 2: NULA with different number of particles
@@ Appendix A Helpful lemmas
Lemma 2.
Suppose $D_{\mu}F:\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ admits a continuous first variation ${\delta D_{\mu}F}:\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$. Then, $D_{\mu}F$ is $\mathscr{L}$-Lipschitz with respect to $W_{1}$ distance satisfying
$\|D_{\mu}F(\mu,x)-D_{\mu}F(\mu^{\prime},x)\|\leq\mathscr{L}W_{1}(\mu,\mu^{\prime})$ (21)
with $\mathscr{L}\triangleq\sup_{\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})}\sup_{x,x^{\prime}\in\mathbb{R}^{d}}\left\|{\delta D_{\mu}F}(\mu,x,x^{\prime})\right\|_{\textsf{\em op}}$
Proof.
By the definition of functional derivative, we have
$\|D_{\mu}F(\mu,x)-D_{\mu}F(\mu^{\prime},x)\|\leq\int_{0}^{1}\left\|\int{\delta D_{\mu}F}((1-t)\mu+t\mu^{\prime},x,x^{\prime})(\mu-\mu^{\prime})\mathrm{d}x^{\prime}\right\|\mathrm{d}t$ (22)
By Kantorovich duality, we obtain
$\left\|\int{\delta D_{\mu}F}((1-t)\mu+t\mu^{\prime},x,x^{\prime})(\mu-\mu^{\prime})\mathrm{d}x^{\prime}\right\|\leq\mathscr{L}W_{1}(\mu,\mu^{\prime}).$
Combining with (22), we complete the proof. ∎
@@
$\leq h\|v_{kh}\|+\left\|\int_{kh}^{t}\int_{0}^{\tau}\gamma v_{\tau^{\prime}}\mathrm{d}\tau^{\prime}\mathrm{d}\tau\right\|+\left\|\int_{kh}^{t}\int_{kh}^{\tau}{D}_{\mu}F(\mu_{\tau^{\prime}}^{x},x_{\tau^{\prime}})\mathrm{d}\tau^{\prime}\mathrm{d}\tau\right\|+\left\|\int_{kh}^{t}\int_{kh}^{\tau}\sqrt{2\gamma}\mathrm{dB}_{\tau^{\prime}}\mathrm{d}\tau\right\|$
@@
$\quad+\left\|\int_{kh}^{t}\int_{kh}^{\tau}\sqrt{2\gamma}\mathrm{dB}_{\tau^{\prime}}\mathrm{d}\tau\right\|$
@@
$\|v_{t}-v_{kh}\|$ $=\left\|\int_{kh}^{t}\gamma v_{\tau}\mathrm{d}\tau\right\|+\left\|\int_{kh}^{t}{D}_{\mu}F(\mu_{\tau}^{x},x_{\tau})\mathrm{d}\tau\right\|+\left\|\int_{kh}^{t}\sqrt{2\gamma}\mathrm{dB}_{t}\right\|$
@@
$|\text{Ent}(\mu_{0})|$ $=\left|\int\mu_{0}\log\mu_{0}\right|$
$=\frac{d}{2}\log(2\pi)+\frac{1}{2}\mathbb{E}_{\mu_{0}}\|\cdot\|^{2}\lesssim d.$
@@
$\mathbb{E}\|(x_{t},v_{t})\|^{2}$ $\leq\frac{2}{\mathscr{C}_{\textsf{LSI}}}\mathcal{F}(\mu_{0})+2\textbf{m}_{2}^{2}$
$\lesssim\frac{\mathscr{L}d}{\mathscr{C}_{\textsf{LSI}}}+d$
@@
$\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}\|(x^{i}_{t},v^{i}_{t})\|^{2}$ $=\frac{1}{N}\sum_{i=1}^{N}W_{2}^{2}(\mu^{i}_{t},\delta_{0})$
$\leq\frac{2}{N}\sum_{i=1}^{N}W_{2}^{2}(\mu^{i}_{t},\mu_{*})+2W_{2}^{2}(\mu_{*},\delta_{0})$
@@
The second inequality follows from Talagrand’s inequality which can be implied by Assumption 2.3. The third inequality follows from Lemma 5. The fourth inequality follows from Lemma 4 and the last inequality follows that $\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{F}^{N}(\mu^{N}_{t})<0$ along the NULD (Proof of Theorem 2.2 in Chen et al. (2023)) and $\mathcal{F}(\mu_{*})\geq 0$. By the definition of $\mathcal{F}^{N}(\mu^{N})$, we have $\mathcal{F}^{N}(\mu_{0}^{N})=\int(NF(\mu_{\textbf{x}})+\frac{1}{2}\|\textbf{v}\|^{2})\mu_{0}^{N}(\mathrm{d}\textbf{x}\mathrm{d}\textbf{v})+\text{Ent}(\mu_{0}^{N})$. Similar to the proof of Lemma 8, since $(\textbf{x},\textbf{v})\sim\mathcal{N}(0,I_{2Nd})$, we have $\int\frac{1}{2}\|\textbf{v}\|^{2}\mu_{0}^{N}(\mathrm{d}\textbf{x}\mathrm{d}\textbf{v})\lesssim Nd$ and $|\text{Ent}(\mu_{0}^{N})|\lesssim Nd$. By Assumption 2.7 and Assumption 2.5, we also have $\int NF(\mu_{\textbf{x}})\mu_{0}^{N}(\mathrm{d}\textbf{x}\mathrm{d}\textbf{v})\lesssim N\mathscr{L}d$ and $\textbf{m}_{2}^{2}\lesssim d$. Thus we have
@@
$\mathrm{d}x_{t}$ $=b_{t}^{\textbf{P}}\mathrm{d}t+\sigma\mathrm{dB}^{\textbf{P}}_{t}\quad\text{under }\textbf{P}_{T},$
$\mathrm{d}x_{t}$ $=b_{t}^{\textbf{Q}}\mathrm{d}t+\sigma\mathrm{dB}^{\textbf{Q}}_{t}\quad\text{under }\textbf{Q}_{T},$
@@
$\frac{\mathrm{d}\mathbf{P}_{T}}{\mathrm{d}\mathbf{Q}_{T}}=\exp\left(\int_{0}^{T}\langle\sigma_{t}^{{\dagger}}(b_{t}^{\textbf{P}_{T}}-b_{t}^{\textbf{Q}_{T}}),\mathrm{dB}^{\textbf{Q}_{T}}_{t}\rangle-\frac{1}{2}\int_{0}^{T}\|\sigma_{t}^{{\dagger}}(b_{t}^{\textbf{P}_{T}}-b_{t}^{\textbf{Q}_{T}})\|^{2}\mathrm{d}t\right)$
Besides, $(\tilde{\mathrm{B}}_{t})_{t\in[0,T]}$ defined by $\mathrm{d}\tilde{\mathrm{B}}_{t}\triangleq\mathrm{dB}_{t}+\sigma_{t}^{{\dagger}}(b_{t}^{Y}-b_{t}^{X})$ is a $\mathbf{P}_{T}$-Brownian motion.
@@ Appendix B Verification of assumptions
$W_{1}$ smoothness has been verified for training mean-field neural networks in Chen et al. (2022). Thus we only verify $W_{1}$ smoothness for examples of density estimation via MMD minimization and KSD minimization. Lemma 2 provides sufficient conditions for $W_{1}$ smoothness. In particular, we have
$\|D_{\mu}F(\mu,x)-D_{\mu}F(\mu^{\prime},x^{\prime})\|\leq\|D_{\mu}F(\mu,x)-D_{\mu}F(\mu^{\prime},x)\|+\|D_{\mu}F(\mu^{\prime},x)-D_{\mu}F(\mu^{\prime},x^{\prime})\|$ (28)
Suzuki et al. (2023) verify that $\|D_{\mu}F(\mu^{\prime},x)-D_{\mu}F(\mu^{\prime},x^{\prime})\|\leq\mathscr{L}\|x-x^{\prime}\|$ for three examples mentioned above. Thus it suffices to verify (21) for the last two examples.
MMD minimization
According to the definition of ^$\hat{\mathcal{M}}(\mu)$ in Section 4, the intrinsic derivative of $F(\mu)$ is
$D_{\mu}F(\mu,x)$ $=D_{\mu}\hat{\mathcal{M}}(\mu,x)+\frac{\lambda^{\prime}}{2}\|x\|^{2}$
$={2\iiint\nabla_{x}p(x;z)p(x^{\prime};z^{\prime})k(z,z^{\prime})\mathrm{d}z\mathrm{d}z^{\prime}\mathrm{d}\mu(x^{\prime})}-{\frac{2}{n}\sum_{i=1}^{n}\int\nabla_{x}p(x;z)k(z,z_{i})\mathrm{d}z}+\frac{\lambda^{\prime}}{2}\|x\|^{2}$
We only need to prove $D_{\mu}\hat{\mathcal{M}}(\mu,x)$ is smooth. The first variation of $D_{\mu}\hat{\mathcal{M}}(\mu,x)$ is
${\delta D_{\mu}\hat{\mathcal{M}}}(\mu,x,x^{\prime})$ $=2\iint\nabla_{x}p(x;z)p(x^{\prime};z^{\prime})k(z,z^{\prime})\mathrm{d}z\mathrm{d}z^{\prime}$
$=-\frac{2}{(2\pi\sigma^{2})^{d}\sigma^{2}}\iint(x-z)\exp\left(-\frac{\|x-z\|^{2}}{2\sigma^{2}}\right)\exp\left(-\frac{\|x^{\prime}-z^{\prime}\|^{2}+\|z-z^{\prime}\|^{2}}{2\sigma^{2}}\right)\mathrm{d}z\mathrm{d}z^{\prime}$
Thus we have
$\left\|{\delta D_{\mu}\hat{\mathcal{M}}}(\mu,x,x^{\prime})\right\|$ $\leq\frac{2}{(2\pi\sigma^{2})^{d}\sigma^{2}}\iint\|x-z\|\exp\left(-\frac{\|x-z\|^{2}}{2\sigma^{2}}\right)\exp\left(-\frac{\|x^{\prime}-z^{\prime}\|^{2}+\|z-z^{\prime}\|^{2}}{2\sigma^{2}}\right)\mathrm{d}z\mathrm{d}z^{\prime}$
$\leq\frac{2}{(2\pi\sigma^{2})^{d}\sigma}\iint\exp\left(-\frac{\|x^{\prime}-z^{\prime}\|^{2}+\|z-z^{\prime}\|^{2}}{2\sigma^{2}}\right)\mathrm{d}z\mathrm{d}z^{\prime}=\frac{2}{\sigma}$
According to Lemma 2 and (28), $F(\mu)$ defined in (29) satisfies Assumption 2.2.
KSD minimization
We now prove that objective (17) satisfies Assumption 2.2 with kernel
$k(x,x^{\prime})=\exp\left(-\frac{\|x\|^{2}}{2\sigma_{1}^{2}}-\frac{\|x^{\prime}\|^{2}}{2\sigma_{1}^{2}}-\frac{\|x-x^{\prime}\|^{2}}{2\sigma_{2}^{2}}\right)$ (30)
and $r(x)=\|x\|^{2}/2$. We also assume the score function of $\mu_{*}$ satisfies (18). Under this assumption on score function and with this choice of kernel, Suzuki et al. (2023) show in their Appendix A that the Stein kernel $u_{\mu_{*}}$ satisfies $\sup_{x,x^{\prime}\in\mathbb{R}^{d}}\max\{|u_{\mu_{*}}|,\|\nabla_{x}u_{\mu_{*}}\|,\|\nabla_{x}^{2}u_{\mu_{*}}\|_{\textsf{op}}\}\leq\mathscr{L}$. With the specific choice of $r(x)=\|x\|^{2}/2$, we reformulate (17) as
$F(\mu)=\textsf{KSD}(\mu)+\frac{\lambda^{\prime}}{2}\mathbb{E}_{x\sim\mu}\|x\|^{2}.$ (31)
Similarly, we only need to verify that $\textsf{KSD}(\mu)$ is smooth with respect to $W_{1}$ distance. The intrinsic derivative of $\textsf{KSD}(\mu)$ is
$D_{\mu}\textsf{KSD}(\mu,x)=\int\nabla_{x}u_{\mu_{*}}(x,x^{\prime})\mathrm{d}\mu(x^{\prime}).$
The first variation of $D_{\mu}\textsf{KSD}(\mu,x)$ is
$\delta D_{\mu}\textsf{KSD}(\mu,x,x^{\prime})=\nabla_{x}u_{\mu_{*}}(x,x^{\prime})$
The following relation implies Assumption 2.2 by Lemma 2.
$\|\delta D_{\mu}\textsf{KSD}(\mu,x,x^{\prime})\|=\|\nabla_{x}u_{\mu_{*}}(x,x^{\prime})\|\leq\mathscr{L}$
B.2 Verification of Assumption 2.5
@@
KSD minimization
Assume the score function $s_{\mu_{*}}$ satisfies (18) and choose the kernel $k$ to be (30), and the Stein kernel $u_{\mu_{*}}$ satisfies $\sup_{x,x^{\prime}\in\mathbb{R}^{d}}\max\{|u_{\mu_{*}}|,\|\nabla_{x}u_{\mu_{*}}\|,\|\nabla_{x}^{2}u_{\mu_{*}}\|_{\textsf{op}}\}\leq\mathscr{L}$ (Suzuki et al., 2023). We now prove the following objective
$F(\mu)=\textsf{KSD}(\mu)+\frac{\lambda^{\prime}}{2}\mathbb{E}_{x\sim\mu}\|x\|^{2}$ (38)
satisfies Assumption 2.5, with $\textsf{KSD}(\mu)$ defined in (16). The functional derivative of $\textsf{KSD}(\mu)$ is
$\frac{\delta\textsf{KSD}}{\delta\mu}(\mu,x)=\int u_{\mu_{*}}(x,x^{\prime})\mathrm{d}\mu(x^{\prime}).$
The functional derivative is bounded as
$\left|\frac{\delta\textsf{KSD}}{\delta\mu}(\mu,x)\right|\leq\int|u_{\mu_{*}}(x,x^{\prime})|\mathrm{d}\mu(x^{\prime})\leq\mathscr{L}.$
$\mathbb{E}_{\hat{\mu}}\|\cdot\|^{2}$ $=\frac{1}{Z}\int\|x\|^{2}\exp\left(-\frac{\delta\textsf{KSD}}{\delta\mu}(\mu,x)-\frac{\lambda^{\prime}}{2}\|x\|^{2}\right)\mathrm{d}x\triangleq\frac{Z^{\prime}}{Z}$ (39)
Now we bound $Z^{\prime}$ and $Z$ respectively.
Thus we have $\mathbb{E}_{\hat{\mu}}\|\cdot\|^{2}=\frac{Z^{\prime}}{Z}\lesssim\frac{\exp(2\mathscr{L})d\lambda^{\prime\frac{d}{2}-1}}{(2\pi)^{\frac{d}{2}}}\lesssim d$ for $\lambda^{\prime}\lesssim(2\pi)^{\frac{d}{d-2}}\exp\left(-\frac{4\mathscr{L}}{d-2}\right)$.
B.3 Verification of Assumption 2.6
@@
  • If $|h(x;a)|\leq\sqrt{\mathscr{L}}(1+\|x\|)$, we have $|h(\mu_{0};a)|\leq\sqrt{\mathscr{L}}\int(1+\|x\|)\mu_{0}(\mathrm{d}x)\lesssim\sqrt{\mathscr{L}}d^{1/2}$. If $l$ is $\sqrt{\mathscr{L}}$-Lipschitz, we have $|l(h(\mu_{0};a_{i}),b_{i})|\leq\sqrt{\mathscr{L}}|h(\mu_{0};a_{i})-b_{i}|\lesssim\mathscr{L}d^{1/2}+\sqrt{\mathscr{L}}\max_{i=1}^{n}|b_{i}|$. We can normalize the data samples to ensure $\max_{i=1}^{n}|b_{i}|\lesssim d\land\sqrt{\mathscr{L}}$. Thus we have $F(\mu_{0})=O(\mathscr{L}d+d)$.
MMD minimization
@@
KSD minimization
Consider the same objective in (17) with $r(x)=\|x\|^{2}/2$ and $\mathcal{N}(0,I_{d})$:
$F(\mu)=\textsf{KSD}(\mu)+\frac{\lambda^{\prime}}{2}\mathbb{E}_{x\sim\mu}\|x\|^{2}.$
If we choose kernel $k(x,x^{\prime})=\exp\left(-\frac{\|x\|^{2}}{2\sigma_{1}^{2}}-\frac{\|x^{\prime}\|^{2}}{2\sigma_{1}^{2}}-\frac{\|x-x^{\prime}\|^{2}}{2\sigma_{2}^{2}}\right)$ and assume the score function of $\mu_{*}$ satisfies $\max\{\|\nabla\log\mu_{*}(x)\|,\|\nabla^{\otimes 2}\log\mu_{*}(x)\|_{\textsf{op}},\|\nabla^{\otimes 3}\log\mu_{*}(x)\|_{\textsf{op}}\}\leq\mathscr{L}(1+\|x\|)$, then the Stein kernel $u_{\mu_{*}}$ satisfies $\sup_{x,x^{\prime}\in\mathbb{R}^{d}}\max\{|u_{\mu_{*}}|,\|\nabla_{x}u_{\mu_{*}}\|,\|\nabla_{x}^{2}u_{\mu_{*}}\|_{\textsf{op}}\}\leq\mathscr{L}$ according to the statement of Appendix A in Suzuki et al. (2023). We have
$F(\mu_{0})$ $=\textsf{KSD}(\mu_{0})+\frac{\lambda^{\prime}}{2}\mathbb{E}_{x\sim\mu_{0}}\|x\|^{2}$
$=\iint u_{\mu_{*}}(x,x^{\prime})\mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime})+\frac{\lambda^{\prime}}{2}\mathbb{E}_{x\sim\mu_{0}}\|x\|^{2}$
$\lesssim\mathscr{L}+d,$
which satisfies Assumption 2.6.
B.4 Verification of Assumption 2.7
@@
which satisfies Assumption 2.7.
KSD minimization
Similar to the verification of Assumption 2.6 above, we have the following relation for $\mu_{0}^{N}=\mathcal{N}(0,I_{Nd})$ under the same assumptions on the score function and kernel:
$\mathbb{E}_{\textbf{x}_{0}\sim\mu_{0}}F(\mu_{\textbf{x}_{0}})$ $=\textsf{KSD}(\mu_{\textbf{x}_{0}})+\frac{\lambda^{\prime}}{2}\mathbb{E}_{\textbf{x}_{0}\sim\mu_{0}^{N}}\frac{1}{N}\sum_{s=1}^{N}\left[\|x_{0}^{s}\|^{2}\right]$
$=\mathbb{E}_{\textbf{x}_{0}\sim\mu_{0}}\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}u_{\mu_{*}}(x_{0}^{i},x_{0}^{j})+\frac{\lambda^{\prime}}{2}\mathbb{E}_{\textbf{x}_{0}\sim\mu_{0}^{N}}\frac{1}{N}\sum_{s=1}^{N}\left[\|x_{0}^{s}\|^{2}\right]\lesssim\mathscr{L}+d,$